find least number of 4 digit which is exactlly divisible by 15, 25, 30?

Answer by Edwin McCravy(20060)

First find the least common multiple:

15 =   3·5
25 =     5²
30 = 2·3 5

So the LCM must contain factors of
one 2, one 3, and two 5's

LCM = 2·3·5² = 150

Then we see what the smallest whole
number that we can multiply 150 by
to get a number larger than 1000, the
smallest 4-digit number. So we divide
1000 by 150 and get 6.66666, so the
smallest whole number we can multiply
150 by to get a 4 digit number is the
smallest whole number greater than
6.66666 which is 7, so the answer is
7·150 or 1050.

Edwin